OpenStudy (anonymous):

help medal for answer and owl bucks :) Solve the equation the square root of the quantity x plus 4 minus 3 equals 1 for the variable. Show each step of your solution process.

OpenStudy (phi):

can you post this using the equation editor (reading math in words is tough)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\sqrt{x+4}-3=1\]

OpenStudy (phi):

can you add 3 to both sides? write + 3 on both sides then simplify what do you get ?

OpenStudy (anonymous):

x=21

OpenStudy (phi):

don't skip steps write +3 on each side, like this \[ \sqrt{x+4}-3+3=1+3 \]

OpenStudy (phi):

on the right side you have 1+3 that simplifies on the left side notice you have -3 + 3 and that simplifies what do you get after simplifying?

OpenStudy (phi):

what is 1+ 3?

OpenStudy (anonymous):

4

OpenStudy (phi):

and what is -3+3 ?

OpenStudy (phi):

-3+3 is the same as 3 + -3 or 3-3